# A Survey for Foundation Models in Autonomous Driving

### v3

@@ 1 Introduction
Despite considerable research in applying foundation models to AD, there are notable limitations and gaps in real-world application. Our survey aims to provide a systematic review and propose future research directions. There are two surveys related to foundation models for autonomous driving: LLM4DriveYang et al. (2023c) is more focused on large language models. Huang et al. (2023) has a good breadth of summary of applications of foundation models in autonomous driving, mainly in simulation, data annotation, and planning. We expand upon existing surveys by covering vision foundation models and multi-modal foundation models, analyzing their applications in prediction and perception tasks. This comprehensive approach includes detailed examinations of technical aspects, such as pre-trained models and methods, and identifies future research opportunities. Innovatively, we propose a taxonomy categorizing foundation models in AD based on modalities and functions, as shown in Figure 1. The following sections will explore the application of various foundation models, including large language models, vision foundation models, and multi-modal foundation models, in the context of AD.
@@ 2.1 Overview
LLMs, originally transformative in NLP, are now driving innovations in AD. Bidirectional Transformers (BERT) Devlin et al. (2018) pioneered foundation models in NLP, leveraging Transformer architecture for understanding language semantics. This pre-trained model can be fine-tuned on specific data-sets, and achieve state-of-the-art results in a wide range of tasks. Following this, OpenAI’s generative pre-trained transformer (GPT) series Radford et al. (2018), including GPT-4, demonstrated remarkable NLP capabilities, attributed to training on extensive datasets. Later GPT models, including ChatGPT, GPT-4Achiam et al. (2023) are trained using billions of parameters and crawled web data with trillions of words, and achieve strong performance on many NLP tasks, including translation, text summarization, question-answering. It also demonstrates one-shot and few-shot reasoning capabilities to learn new skills from the context. More and more researchers have started to apply these reasoning, understanding, and in-context learning capabilities to address challenges in AD.
@@ 2.2.1 Reasoning and Planning
The decision-making process in AD closely parallels human reasoning, necessitating the interpretation of environmental cues to make safe and comfortable driving decisions. LLMs, through their training on diverse web data, have assimilated common-sense knowledge pertinent to driving, drawing from a plethora of sources including web forums and official government websites. This wealth of information enables LLMs to engage in the nuanced decision-making required for AD. One method for harnessing LLMs in AD involves presenting them with detailed textual descriptions of the driving environment, prompting them to propose driving decisions or control commands. This process, as illustrated in Figure 2, typically encompasses comprehensive prompts detailing agent states, such as coordinates, speed, and past trajectories, the vehicle’s state i.e, velocity, and acceleration, and map specifics including traffic lights, lane information, and intended route). For enhanced interaction understanding, LLMs can also be directed to provide reasoning along with their responses. For instance, the GPT driver Mao et al. (2023a) not only recommends vehicle actions but also elucidates the rationale behind these suggestions, significantly enhancing the transparency and explainability of autonomous driving decisions. This approach, exemplified by Driving with LLMs Chen et al. (2023), enhances the explainability of autonomous driving decisions. Similarly, the “Receive, Reason, and React” approach Cui et al. (2023) instructs LLM agents to assess lane occupancy and evaluate the safety of potential actions, thereby fostering a deeper comprehension of dynamic driving scenarios. These methods not only leverage LLMs’ inherent ability to understand complex scenarios but also employ their reasoning capabilities to simulate human-like decision-making processes. Through the integration of detailed environmental descriptions and strategic prompts, LLMs contribute significantly to the planning and reasoning aspects of AD, offering insights and decisions that mirror human judgment and expertise.
@@ 2.2.2 Prediction
Prediction forecasts traffic participants’ future trajectories, intents, and possible interactions with the ego vehicle. The common deep learning-based models are based on rasterized or vector images of the traffic scene, which encode spatial information. However, it is still challenging to accurately predict highly interactive scenes, which requires reasoning and semantic information, for example, right-of-ways, vehicles’ turning signals, and pedestrians’ gestures. The text representation of the scene can provide more semantic context and better leverage LLM’s reasoning capability and common knowledge in the pre-training dataset.
Keysan et al. (2023) did an early exploration of LLM’s power to make trajectory predictions. They convert the scene representation into text prompts, and use BERTDevlin et al. (2018) model to generate the text encoding, which is finally fused with image encoding to decode trajectory prediction. Their evaluation shows significant improvement compared with baselines only using image encoding or text encoding.
LG-TrajChib and Singh (2024) leverages pre-trained LLM to generate motion cues for pedestrian trajectory prediction. Given the prompt and pedestrian coordinates, LLM is able to identify pedestrian motion patterns, such as linear motion, curved motion, and standing still. Combined with the Transformer-based motion encoder and social decoder, their method achieved the best results in multiple data sets compared with other state-of-the-art methods.
iMotion-LLMFelemban et al. (2024) proposes a novel motion prediction method leveraging instructions and captions in text. Given the text instruction inputs and scene encoding projected into text space, LLM decodes the output captions describing the driving decisions and latent tokens to generate multi-modal trajectories. Their model significantly outperforms the baseline model without using LLM. Additionally, they propose a feasibility classification task, which aligns predicted trajectories with feasible instructions and reject infeasible ones to improve safety.
@@ 2.2.3 User Interface and Personalization
Autonomous vehicles should be user-friendly and able to follow the instructions from passengers or remote operators. The current Robotaxi remote assistance interface is only used to execute a limited set of pre-defined commands. However, LLM’s understanding and interaction capabilities make it possible to let autonomous driving cars to understand human’s free-form instruction to better control the autonomous vehicle and satisfy users’ personalized requirements. Cui et al. (2023) explores the LLM-based planner conditioning on personalized commands, e.g. “driving aggressively” or “conservatively”, and was able to output actions of various speeds and riskiness. Yang et al. (2023b) leverages LLM’s reasoning abilities, and provides step-by-step rules to decide on response to user commands. The LLM agent is also able to accept or reject user commands based on pre-defined traffic rules and system requirements.
@@ 2.2.4 Simulation and Testing
LLM can summarize and extract knowledge from existing text data and generate new content, which can facilitate simulation and testing. The ADEPT systemWang et al. (2022b) uses GPT to extract key information from NHTSA accident reports using QA approach, and was able to generate diverse scene code used for simulation and testing. TARGETDeng et al. (2023) system is able to use GPT to translate traffic rules from the natural language to the domain-specific language, which is used for generating testing scenarios. LCTGenTan et al. (2023) uses LLM as a powerful interpreter translating user’s text query into structured specifications of map lanes and vehicle locations for traffic simulation scenarios.
@@ 2.3 Methods and Techniques
Researchers use similar techniques in natural language processing to utilize LLM for autonomous driving tasks, such as prompt engineering, in-context and few-shot learning, and reinforcement learning from human feedbackOuyang et al. (2022).
@@ 2.3.1 Prompt Engineering
Some papers add traffic rules as pre-prompt to make LLM agent law-compliant. Driving with LLMsChen et al. (2023) has diving rules covering aspects like traffic light transition and left or right driving side. Mao et al. (2023b) proposes a module called common-sense module, which stores the rules and instructions for human driving, for example, avoiding collision, and maintaining safety distances.
LanguageMPCSha et al. (2023) adopts a top-down decision-making system: given different situations, the vehicle has different possible actions. LLM agent is also instructed to identify important agents in the scenario and output attention, weight, and bias matrices to select from pre-defined actions.
Memory modules are also introduced in some papers, which store past driving scenarios. At inference time, the relevant examples are retrieved and added as the context in the prompt, and LLM agent can better leverage few-shot learning capabilities and reflect on the most relevant scenarios. DILUWen et al. (2023a) proposes a memory module, which stores text descriptions of driving scenarios in the vector database, and the system can retrieve top-k scenarios for few-shot learning. Mao et al. (2023b) has a two-stage retrieval process: the first stage uses k-nearest-neighbor search to retrieve relevant past examples in the database, and the second stage asks LLM to rank these examples.
More papers built complex systems to manage tasks in the prompt generation, which trigger function calls to other modules or sub-systems to obtain required information for decision-making. Mao et al. (2023b) has created libraries and function API calls to interact with perception, prediction, and mapping systems so that the LLM can fully leverage all available information. LanguageMPC Sha et al. (2023) uses LangChain to create tools and interfaces needed by LLM to get relevant vehicles, possible situations, and available actions.
@@
Fine-tuning and in-context learning are both applied to adapt pre-trained models to autonomous driving. Fine-tuning re-trains the model parameters on smaller domain-specific datasets, while in-context learning or few-shot learning leverages LLM’s knowledge and reasoning ability to learn from given examples in the input prompt. Most papers are focused on in-context learning, but only a few papers utilize fine-tuning. Researchers have mixed results on which one is the better: Mao et al. (2023b) compared both approaches and found that few-shot learning is slightly more effective. GPT-Driver Mao et al. (2023a) has a different conclusion that using OpenAI fine-tuning performs significantly better than few-shot learning. Chen et al. (2023) also compared training from scratch and fine-tuning approaches, and found using the pre-trained LLaMA model with LoRA-based fine-tuning can perform better than training from scratch.
@@ 2.3.3 Reinforcement Learning and Human Feedback
DILU Wen et al. (2023a) proposes reflection modules, which store good driving examples and bad driving examples with human corrections to enhance its reasoning capabilities further. In this way, the LLM can learn to reason about what action is safe and unsafe and continuously reflect on a large amount of past driving experiences. Surreal Driver Jin et al. (2023) interviewed 24 drivers and used their descriptions of driving behavior as chain-of-thought prompts to develop a ‘coach agent’ module, which can instruct the LLM model to have a human-like driving style. Incorporating Voice Instructions Wang et al. (2022a) uses instructions from human coaches, and builds a taxonomy of natural language instructions of action, reward, and reasoning, which are used to train a deep reinforcement learning-based autonomous driving agent.
@@ 2.4.1 Hallucination and Harmfulness
Hallucination is a big challenge in LLM, and state-of-the-art large language models can still produce misleading and false information. Most methods proposed in existing papers still require parsing driving actions from LLM’s response. When given an unseen scenario, the LLM model can still produce unhelpful or wrong driving decisions. Autonomous driving is a safety-critical application, which has much higher reliability and safety requirements than chat-bots. According to evaluation result Mao et al. (2023a), the LLM model for autonomous driving has a 0.44% collision rate, higher than other methods. Chen et al. (2023) proposes a method to reduce hallucination by asking questions without enough information to make decisions, and instructs LLM to answer “I don’t know”. The pre-trained LLM may also include harmful content, for example, aggressive driving and speeding. More human-in-the-loop training and alignment can reduce hallucinations and harmful driving decisions.
@@ 2.4.2 Latency and Efficiency
Large language models often suffer from high latency, and generating detailed driving decisions can exhaust the latency budget of limited compute resources in the car. It takes several seconds for inference according to Jin et al. (2023). LLMs with billions of parameters can consume over 100GB of memory, which might interfere with other critical modules in autonomous driving vehicles. More research needs to be done in this domain, such as model compression and knowledge distillation, to make LLM more efficient and easier for deployment.
@@ 2.4.3 Dependency on Perception System
Despite the supreme reasoning capability of LLM, the environment description still depends on the upstream perception module. The driving decisions can go wrong and cause critical accidents with minor errors in environmental inputs. For example, Mao et al. (2023b) shows failure cases when upstream heading data has errors. LLM also needs to be better adapted to perception models and make better decisions when there are errors and uncertainty.
@@ 3 Vision Foundation Model
Vision foundation models have achieved great success in multiple computer vision tasks, such as object detection and segmentation. DINO Caron et al. (2021) uses vision-transformer architecture, and is trained in a self-supervised manner, predicting global image features given local image patches. DINOV2Oquab et al. (2023) scales the training with one billion parameters and a diversely curated dataset of 1.2 billion images and achieves state-of-the-art results in multiple tasks. Segment-anything modelKirillov et al. (2023) is a foundation model for image-segmentation. The model is trained with different types of prompts (points, boxes, or texts) to generate segmentation masks. Trained with billions of segmentation masks in the dataset, the model shows zero-shot transfer capability to segment new objects given the appropriate prompt.
Diffusion modelSohl-Dickstein et al. (2015) is a generative foundation model widely used for image generation. The diffusion model iteratively adds noise to the image and applies a reverse diffusion process to restore the image. To generate the image, we can sample from the learned distribution and restore highly realistic images from random noises. Stable-DiffusionRombach et al. (2022) model uses VAEKingma and Welling (2013) to encode images to latent representation and use UNetRonneberger et al. (2015) to decode from latent variable to pixel-wise images. It also has an optional text encoder and applies the cross-attention mechanism to generate images conditional on prompts (text description or other images). DALL-ERamesh et al. (2021) model was trained with billions of image and text pairs and uses stable diffusion to generate high-fidelity images and creative arts following human instructions.
There is growing interest in the application of vision foundation models in autonomous driving, mainly for 3D perception and video generation tasks.

### v4

@@ 1 Introduction
Despite considerable research in applying foundation models to AD, there are notable limitations and gaps in real-world application. Our survey aims to provide a systematic review and propose future research directions. There are two surveys related to foundation models for autonomous driving: LLM4Drive [1] is more focused on large language models. [2] has a good breadth of summary of applications of foundation models in autonomous driving, mainly in simulation, data annotation, and planning. We expand upon existing surveys by covering vision foundation models and multi-modal foundation models, analyzing their applications in prediction and perception tasks. This comprehensive approach includes detailed examinations of technical aspects, such as pre-trained models and methods, and identifies future research opportunities. Innovatively, we propose a taxonomy categorizing foundation models in AD based on modalities and functions, as shown in Figure 1. In the following sections, we will explore the application of various foundation models, including large language models, vision foundation models, and multi-modal foundation models, in the context of AD.
@@ 2.1 Overview
LLMs, originally transformative in NLP, are now driving innovations in AD. Bidirectional Transformers (BERT) [3] pioneered foundation models in NLP, leveraging Transformer architecture for understanding language semantics. This pre-trained model can be fine-tuned on specific data-sets, and achieve state-of-the-art results in a wide range of tasks. Following this, OpenAI’s generative pre-trained transformer (GPT) series [4], including GPT-4, demonstrated remarkable NLP capabilities, attributed to training on extensive datasets. Later GPT models, including ChatGPT, GPT-4 [5] are trained using billions of parameters and crawled web data with trillions of words, and achieve strong performance on many NLP tasks, including translation, text summarization, question-answering. It also demonstrates one-shot and few-shot reasoning capabilities to learn new skills from the context. More and more researchers have started to apply these reasoning, understanding, and in-context learning capabilities to address challenges in AD.
@@ 2.2.1 Reasoning and Planning
The decision-making process in AD closely parallels human reasoning, necessitating the interpretation of environmental cues to make safe and comfortable driving decisions. LLMs, through their training on diverse web data, have assimilated common-sense knowledge pertinent to driving, drawing from a plethora of sources including web forums and official government websites. This wealth of information enables LLMs to engage in the nuanced decision-making required for AD. One method for harnessing LLMs in AD involves presenting them with detailed textual descriptions of the driving environment, prompting them to propose driving decisions or control commands. This process, as illustrated in Figure 2, typically encompasses comprehensive prompts detailing agent states, such as coordinates, speed, and past trajectories, the vehicle’s state i.e, velocity, and acceleration, and map specifics including traffic lights, lane information, and intended route). For enhanced interaction understanding, LLMs can also be directed to provide reasoning along with their responses. For instance, the GPT driver [6] not only recommends vehicle actions but also elucidates the rationale behind these suggestions, significantly enhancing the transparency and explainability of autonomous driving decisions. This approach, exemplified by Driving with LLMs [7], enhances the explainability of autonomous driving decisions. Similarly, the “Receive, Reason, and React” approach [8] instructs LLM agents to assess lane occupancy and evaluate the safety of potential actions, thereby fostering a deeper comprehension of dynamic driving scenarios. These methods not only leverage LLMs’ inherent ability to understand complex scenarios but also employ their reasoning capabilities to simulate human-like decision-making processes. Through the integration of detailed environmental descriptions and strategic prompts, LLMs contribute significantly to the planning and reasoning aspects of AD, offering insights and decisions that mirror human judgment and expertise.
@@ 2.2.2 Prediction
Prediction forecasts traffic participants’ future trajectories, intents, and possible interactions with the ego vehicle. The common deep learning-based models are based on rasterized or vector images of the traffic scene, which encode spatial information. However, it is still challenging to accurately predict highly interactive scenes, which requires reasoning and semantic information, for example, right-of-ways, vehicles’ turning signals, and pedestrians’ gestures. The text representation of the scene can provide more semantic information, and better leverage LLM’s reasoning capability and common knowledge in the pre-training dataset. There is still not much research applying LLM to trajectory prediction. [9] did an early exploration of LLM’s power to make trajectory predictions. They convert the scene representation into text prompts, and use BERT model to generate the text encoding, which is finally fused with image encoding to decode trajectory prediction. Their evaluation shows significant improvement compared with baselines only using image encoding or text encoding.
@@ 2.2.3 User Interface and Personalization
Autonomous vehicles should be user-friendly and able to follow the instructions from passengers or remote operators. The current Robotaxi remote assistance interface is only used to execute a limited set of pre-defined commands. However, LLM’s understanding and interaction capabilities make it possible to let autonomous driving cars to understand human’s free-form instruction to better control the autonomous vehicle and satisfy users’ personalized requirements. [8] explores the LLM-based planner conditioning on personalized commands, e.g. “driving aggressively” or “conservatively”, and was able to output actions of various speeds and riskiness. [10] leverages LLM’s reasoning abilities, and provides step-by-step rules to decide on response to user commands. The LLM agent is also able to accept or reject user commands based on pre-defined traffic rules and system requirements.
@@ 2.2.4 Simulation and Testing
LLM can summarize and extract knowledge from existing text data and generate new content, which can facilitate simulation and testing. The ADEPT system [11] uses GPT to extract key information from NHTSA accident reports using QA approach, and was able to generate diverse scene code used for simulation and testing. TARGET [12] system is able to use GPT to translate traffic rules from the natural language to the domain-specific language, which is used for generating testing scenarios. LCTGen [13] uses LLM as a powerful interpreter translating user’s text query into structured specifications of map lanes and vehicle locations for traffic simulation scenarios.
@@ 2.3 Methods and Techniques
Researchers use similar techniques in natural language processing to utilize LLM for autonomous driving tasks, such as prompt engineering, in-context and few-shot learning, and reinforcement learning from human feedback (RLHF)[14].
@@ 2.3.1 Prompt Engineering
Some papers add traffic rules as pre-prompt to make LLM agent law-compliant. Driving with LLMs [7] has diving rules covering aspects like traffic light transition and left or right driving side. [15] proposes a module called common-sense module, which stores the rules and instructions for human driving, for example, avoiding collision, and maintaining safety distances.
LanguageMPC [16] adopts a top-down decision-making system: given different situations, the vehicle has different possible actions. LLM agent is also instructed to identify important agents in the scenario and output attention, weight, and bias matrices to select from pre-defined actions.
Memory modules are also introduced in some papers, which store past driving scenarios. At inference time, the relevant examples are retrieved and added as the context in the prompt, and LLM agent can better leverage few-shot learning capabilities and reflect on the most relevant scenarios. DILU[17] proposes a memory module, which stores text descriptions of driving scenarios in the vector database, and the system can retrieve top-k scenarios for few-shot learning. [15] has a two-stage retrieval process: the first stage uses k-nearest-neighbor search to retrieve relevant past examples in the database, and the second stage asks LLM to rank these examples.
More papers built complex systems to manage tasks in the prompt generation, which trigger function calls to other modules or sub-systems to obtain required information for decision-making. [15] has created libraries and function API calls to interact with perception, prediction, and mapping systems so that the LLM can fully leverage all available information. LanguageMPC [16] uses LangChain to create tools and interfaces needed by LLM to get relevant vehicles, possible situations, and available actions.
@@
Fine-tuning and in-context learning are both applied to adapt pre-trained models to autonomous driving. Fine-tuning re-trains the model parameters on smaller domain-specific datasets, while in-context learning or few-shot learning leverages LLM’s knowledge and reasoning ability to learn from given examples in the input prompt. Most papers are focused on in-context learning, but only a few papers utilize fine-tuning. Researchers have mixed results on which one is the better: [15] compared both approaches and found that few-shot learning is slightly more effective. GPT-Driver [6] has a different conclusion that using OpenAI fine-tuning performs significantly better than few-shot learning. [7] also compared training from scratch and fine-tuning approaches, and found using the pre-trained LLaMA model with LoRA-based fine-tuning can perform better than training from scratch.
@@ 2.3.3 Reinforcement Learning and Human Feedback
DILU [17] proposes reflection modules, which store good driving examples and bad driving examples with human corrections to enhance its reasoning capabilities further. In this way, the LLM can learn to reason about what action is safe and unsafe and continuously reflect on a large amount of past driving experiences. Surreal Driver [18] interviewed 24 drivers and used their descriptions of driving behavior as chain-of-thought prompts to develop a ‘coach agent’ module, which can instruct the LLM model to have a human-like driving style. Incorporating Voice Instructions [19] uses instructions from human coaches, and builds a taxonomy of natural language instructions of action, reward, and reasoning, which are used to train a deep reinforcement learning-based autonomous driving agent.
@@ 2.4.1 Hallucination and Harmfulness
Hallucination is a big challenge in LLM to avoid, and state-of-the-art large language models can still produce misleading and false information, not only in text-based models but also in scenarios like multimodality-based tasks. Hallucination happens when a model generates output that contains fictional, misleading, or entirely fabricated details, facts, or claims, instead of delivering reliable and truthful information. When talking about its relationship with AD, most methods proposed in existing papers still require parsing driving actions from LLM’s response. When given an unseen scenario, the LLM model can still produce unhelpful or wrong driving decisions. Autonomous driving is a safety-critical application, which has much higher reliability and safety requirements than chat-bots. According to evaluation result [6], the LLM model for autonomous driving has a 0.44% collision rate, higher than other methods. [7] proposes a method to reduce hallucination by asking questions without enough information to make decisions, and instructs LLM to answer “I don’t know”. The pre-trained LLM may also include harmful content, for example, aggressive driving and speeding. More human-in-the-loop training and alignment (like RLHF [20] and DPO [21]) can reduce hallucinations and harmful driving decisions.
@@ 2.4.2 Latency and Efficiency
Large language models often suffer from high latency, and generating detailed driving decisions can exhaust the latency budget of limited compute resources in the car. It takes several seconds for inference according to [18]. LLMs with billions of parameters can consume over 100GB of memory, which might interfere with other critical modules in autonomous driving vehicles. Currently, many methods have been proposed to boost the inference speed for LLM-based decoder models, including model compression and knowledge distillation. A new attention structure like PagedAttention[22] has been proposed for fast inference in the generative model. Quantization methods like GPTQ[23], AWQ[24], SqueezeLLM[25] has been designed to achieve around 2.1x to 2.3x speedup compared with the baseline. Distillation methods like SLidR [26], ST-SLiDR[27] have been proposed for LiDAR style input on detection and segmentation tasks.
@@ 2.4.3 Dependency on Perception System
Despite the supreme reasoning capability of LLM, the environment description still depends on the upstream perception module. The driving decisions can go wrong and cause critical accidents with minor errors in environmental inputs. For example, [15] shows failure cases when upstream heading data has errors. LLM also needs to be better adapted to perception models and make better decisions when there are errors and uncertainty.
@@ 3 Vision Foundation Model
Vision foundation models have achieved great success in multiple computer vision tasks, such as object detection and segmentation. DINO [28] uses vision-transformer architecture, and is trained in a self-supervised manner, predicting global image features given local image patches. DINOV2 [29] scales the training with one billion parameters and a diversely curated dataset of 1.2 billion images and achieves state-of-the-art results in multiple tasks. Segment-anything model [30] is a foundation model for image-segmentation. The model is trained with different types of prompts (points, boxes, or texts) to generate segmentation masks. Trained with billions of segmentation masks in the dataset, the model shows zero-shot transfer capability to segment new objects given the appropriate prompt.
Diffusion model [31] is a generative foundation model widely used for image generation. In recent years, the diffusion model [32] has gained significant acclaim in the field of image synthesis. This model generates images by starting with Gaussian noise and progressively refining it through a series of denoising steps. The methodology is grounded in solid physical principles, encompassing both a diffusion phase and a reverse phase. During the diffusion phase, an image is incrementally transformed into a Gaussian distribution by repeatedly adding random Gaussian noise. Conversely, the reverse phase involves reconstructing the original image from this distribution through multiple denoising iterations. The diffusion model iteratively adds noise to the image and applies a reverse diffusion process to restore the image. To generate the image, we can sample from the learned distribution and restore highly realistic images from random noises. Latent diffusion models (LDM) (also known as the stable-diffusion model) [33] are a type of diffusion model that focuses on the distribution of the latent space of images and have recently demonstrated outstanding performance in image synthesis. An LDM comprises two components: an autoencoder and a diffusion model, as illustrated in Figure 4. The autoencoder is responsible for compressing and reconstructing images, utilizing an encoder and a decoder. The encoder transforms the image into a lower-dimensional latent space, while the decoder reconstructs the original image from this latent space. Subsequently, the latent generative model is trained to replicate a fixed forward Markov chain using DDPMs [34]. Stable-Diffusion [33] model uses VAE [35] to encode images to latent representation and use UNet [36] to decode from latent variable to pixel-wise images. It also has an optional text encoder and applies the cross-attention mechanism to generate images conditional on prompts (text description or other images).
DALL-E [37] model was trained with billions of image and text pairs and uses stable diffusion to generate high-fidelity images and creative arts following human instructions. DALL-E 2 [38], an extension of DALL-E [37], integrates a CLIP encoder with a diffusion decoder to handle both image generation and editing tasks, as depicted in Fig. 19. Unlike Imagen, DALL-E 2 utilizes a prior network to translate between text embeddings and image embeddings. Building on this, DALL-E 3 focuses on enhancing prompt adherence and caption quality. It first trains a robust image captioner capable of generating detailed and accurate image descriptions, and then uses this captioner to produce even more refined and detailed captions. There is growing interest in the application of vision foundation models in autonomous driving, mainly for 3D perception and video generation tasks.
Figure 4: Stable Diffusin